# custom_references.bib

### v1

@@ 7 Conclusion and Open Problems
Question 7.1 (Simultaneous biasing).
@@
Note that a much weaker version of 7.1 indeed holds by Theorem 4.1. The difference is that in Theorem 4.1, the (slightly different) sets of bad players can corrupt each $f_{i}$ differently while here we want each of the functions indexed by $P$ to be simultaneously corrupted not only by the same set of bad players $B$, but also by the same behavior of the bad players in that set.
We note that when $m=O(1)$, by repeatedly applying the result of [KKL88influence], 7.1 does hold. However, that strategy requires more than $n$ players once $m\geq\log\ell$, which is trivial. Answering the above question in the positive will lead to almost matching lower bounds for coin flipping - in particular, this would imply that $\frac{\ell\operatorname{poly}(\log^{(2)}\ell)}{\log\ell}$ bad players suffice to bias any two-round protocol. This bound would follow by using the same strategy as the proof of Theorem 5.1, with the ‘second induction step’ replaced by the simultaneous biasing conjecture. For multi-round lower bounds nearly matching our constructions from Theorem 6.1, we would need a stronger version of 7.1 where we would want to simultaneously bias not just a function map but a $k$-round protocol map.
Acknowledgements
We thank the organizers of the Dagstuhl Seminar on Algebraic and Analytic Methods in Computational Complexity and Schloss Dagstuhl for providing a stimulating research environment, where discussions between R.S. and E.C. initiated this collaboration.
[…]

### v2

@@ 9 Conclusion and Open Problems
Question 9.1 (Simultaneous biasing).
@@
Note that a much weaker version of 9.1 indeed holds by Theorem 4.1. The difference is that in Theorem 4.1, the (slightly different) sets of bad players can corrupt each $f_{i}$ differently while here we want each of the functions indexed by $P$ to be simultaneously corrupted not only by the same set of bad players $B$, but also by the same behavior of the bad players in that set.
We note that when $m=O(1)$, by repeatedly applying the result of [KKL88], 9.1 does hold. However, that strategy requires more than $n$ players once $m\geq\log\ell$, which is trivial. Answering the above question in the positive will lead to almost matching lower bounds for coin flipping; in particular, this would imply that $\frac{\ell\operatorname{poly}(\log^{(2)}\ell)}{\log\ell}$ bad players suffice to bias any two-round protocol. This bound would follow by using the same strategy as the proof of Theorem 5.1, with the second induction step replaced by the simultaneous biasing conjecture. Our one-round random selection impossibility results from 7 can also be interpreted as a necessary step towards this conjecture. For multi-round lower bounds nearly matching our constructions from Theorem 6.1, we would need a stronger version of 9.1 where we would want to simultaneously bias not just a function map but a $k$-round protocol map.
Appendix A Comparison to Proof Strategy of [RSZ02]
Since our coin flipping impossibility results from 1 are inspired by the work of [RSZ02] and our proofs build on their arguments, we here briefly sketch their strategy for 2 rounds and how we improve upon it.
Let $\mathcal{F}=\{\pi_{\alpha}\}$ where $\alpha\in\{0,1\}^{\ell}$ and let $\pi_{\alpha}$ be the induced second round protocol when in the first round the players output $\alpha$. [RSZ02] proceeds by finding a set of bad players in the second round to bias the functions in $\mathcal{F}$. To do this, as hinted at earlier, [RSZ02] analyzes the same semi-random process as in Lemma 2.6. At the end of the process, they also obtain a common random set $B_{R}$ and a specialized heavy set $B_{H}=B_{H}(f)$ such that for most functions $f\in\mathcal{F}$, $B_{R}\cup B_{H}(f)$ can bias the function $f$. They then pick another random set $B_{R^{\prime}}$ and bound the probability that $B_{H}(f)\subset B_{R^{\prime}}$ for fixed $f$, and argue that in expectation, $\delta\approx\left(\frac{r}{\ell}\right)^{2^{\ell/r}}$ fraction of functions $f\in\mathcal{F}$ will be such that $B_{H}(f)\subset B_{R^{\prime}}$ where $\left\lvert B_{R^{\prime}}\right\rvert=r$. They let $B_{R}\cup B_{R^{\prime}}$ be the set of bad players that will bias the second round protocols.
Now, let $S\subset\{0,1\}^{\ell}$ be the set of $\alpha\in\{0,1\}^{\ell}$ such that $\pi_{\alpha}$ can be biased by $B_{R}\cup B_{R^{\prime}}$, and let $g:\{0,1\}^{\ell}\to\{0,1\}$ be such that $g(\alpha)=1\iff\alpha\in S$. We know from above that $\Pr[g=1]\geq\delta$. Now, the goal for the first round is to find bad players to bias $g$ towards $1$ so that with high probability, the resulting second round protocol can be biased using $B_{R}\cup B_{R^{\prime}}$. To do this, [RSZ02] use the KKL theorem [KKL88] which lets them do this by finding $O(\ell/(\delta\log\ell))$ bad players. To minimize the total number of bad players, they end up using $O(\ell/\log^{(3)}\ell)$ bad players. For multiple rounds, their result follows by an inductive argument, showing that $O\left(\ell/\log^{(2k-1)}\ell\right)$ bad players suffice to bias $k$ round protocols.